January 14, 2016 Drift Chamber meeting
- Instructions for Bluejeans meeting connection
- Meeting ID: 290664653
- To join via a Web Browser, go to the page  https://bluejeans.com/290664653.
- NIM Papers
- Spring 2016 Run Planning
- F125 Update (Naomi)
- CDC update (Beni/Mike)
- FDC update (Lubomir)
- Electronics (Fernando, Chris, Nick)
Participants: Naomi, Mike (CMU), Luke, Chris, Nick, Simon, Beni, Sergey and Lubomir (JLab).
- Noami and Simon working on different sections of the CDC paper.
Spring 2016 Run Planning Meeting
- There were no special requirements for the coming run from both CDC and FDC. Naomi and Beni want to do HV scan of the CDC. We plan to run the FDC at a fixed HV. When the beam is stable if the magnet is not yet operational we may take straight track data, but no more than a shift.
- The reason for the inefficiency in the region between the FDC and CDC is not misalignment between the two detectors but tracking issues (Simon). Mike: we see this not only in the data but in the MC, as well.
- Will start long cosmic run over the extended weekend with BCAL trigger reading both CDC and FDC: CDC_short and FDC_long formats.
- Naomi: investigating events with "missing object" and with differences in the fADC-calculated and emulated quantities. Some problems also in the emulations - Dave fixed a lot but not everything.
- Beni also started looking into the "bad" events with his own code.
- No issues, all HV set at 500V, will go to the nominal for the weekend.
- Fixing the bad HV channel requires access to the detector. Should be discussed withing the Hall D and, if needed, we may do this during the summer break.
- Lubomir discussed a plot (see above) from the TRD test with PS trigger showing the x-coordinate (perp. to beam) distribution of the hits on the two chambers and of the reconstructed tracks. The dips for both chambers may come from the PS trigger inefficiency; need to study this with Sasha.